### v2

@@ Abstract
Results. In this paper we present our first results for an inclined dipole rotator using the new methodology. We are able to zoom-in around the Y-point and inside the closed-line region with unprecedented detail, and we observe features that were never been discussed in previous numerical solutions. This is the first time the steady-state 3D problem is addressed directly, and not through a time-dependent simulation that eventually relaxes to a steady-state.
Conclusions. We have trained a Neural Network that instantaneously yields the three components of the magnetic field and their spatial derivatives at any given point. Our results demonstrate the potential of the new method to generate the reference solution of the ideal pulsar magnetosphere.
@@ 1 Gaps in our current understanding
Right after the discovery of pulsars (Hewish, 1968), Goldreich and Julian formulated the mathematical description of their plasma-filled magnetospheres in the ideal force-free limit in steady-state and axisymmetry (Goldreich & Julian, 1969). It took us almost three decades to produce the first solution of this idealized problem, mainly because of the mathematical singularity along the light cylinder (Contopoulos et al., 1999). Since then, this solution became a reference solution against which the results of numerical simulations are evaluated. As we will see below, the ideal force-free steady-state 3D problem is also well formulated mathematically, but still lacks a reference solution. This is mainly due to the presence of electric current sheets within the magnetosphere which strongly complicate all our standard numerical approaches.
@@
Recent advances in global particle-in-cell (PIC) simulations (e.g. Philippov et al., 2015; Hu & Beloborodov, 2022; Hakobyan et al., 2023; Bransgrove et al., 2023; Soudais et al., 2024) have raised several questions. Notably, these simulations yield separatrix regions between open and closed field lines that exhibit a significant thickness beyond the simulation resolution, and a multi-layered internal structure. This is intriguing because, the same simulations are able to generate Harris-type equatorial current sheets with thickness at the limit of their numerical resolution (the skin-depth). Moreover, their closed-line regions terminate well inside the light cylinder, with regions of strong electromagnetic dissipation beyond their tips, something that was not expected by Goldreich and Julian. Interestingly, the shape of the tips resembles a pointed Y, thus deviating from the theoretically predicted T shape (Uzdensky, 2003). Furthermore, the extent of the closed-line region and the sharpness of the Y appears to be linked to the efficiency of pair production in the simulation (see comment in Kalapotharakos et al., 2018). Finally, our recent work in Contopoulos et al. (2024) indicates that as the closed-line region approaches the light cylinder, the electric current in the separatrix decreases, and in the limit where it touches the light cylinder, the equatorial current sheet vanishes entirely.
@@
The origin of these results remains unclear. Some researchers suspect that they arise either due to insufficient resolution in current PIC simulations, or to inadequate running times to reach a true steady-state. We believe that the issue may be more serious because current state-of-the-art numerical methods seem inadequate to treat magnetospheric current sheets, especially the separatrix current sheet that contains a guide field and is connected at its tip to the equatorial current sheet (see the Appendix). On the other hand, the resolution of current PIC simulations falls significantly short in modeling the microphysics of the equatorial current sheet. To generate pulsar light curves and spectra for comparison with observations, simulation results are extrapolated by several orders of magnitude (particle Lorentz factors, magnetic field values, electromagnetic spectra, etc.). Unfortunately, there is no consensus among research groups regarding the specific extrapolation method. Consequently, a universally accepted understanding of the physical origin of high-energy radiation from pulsars remains elusive.
@@
Given the inherent challenges of all traditional numerical methods, we believe that this represents a critical development that could significantly misdirect pulsar research, especially considering that reviews have already begun to emerge in the literature. Our proposal is to revisit fundamental principles and independently derive the reference ideal force-free magnetosphere using a novel numerical approach that treats the separatrix and equatorial discontinuities as genuine ideal current sheets, rather than mere numerical approximations. In the present paper, we generalize in 3D the axisymmetric dissipationless ideal solutions of Dimitropoulos et al. (2024, hereafter Paper I) and Contopoulos et al. (2024, hereafter Paper II). This is the third paper (Paper III) of an ongoing investigation that aims to produce the reference ideal solution of the 3D pulsar magnetosphere against which all past and ongoing numerical simulations will be compared and evaluated.
@@ 2 Domain decomposition
Once the size and shape of the polar cap is determined, the shape of the separatrix surface is unknown and must be self-consistently determined. In the present Letter, we implement in 3D the procedure first proposed in Paper I. We begin by separating the regions of open and closed field lines with an ad hoc chosen surface, namely
@@
(see also Lyubarskii 1990). Treating such discontinuities inside any computational domain is very problematic and all computational methods generate spurious Gibbs oscillations around the separatrix (see discussion in Cao & Yang, 2022). We propose here a better way to treat such discontinuities as follows:
@@
We have implemented one more trick that greatly simplifies our problem, namely a novel numerical treatment of the equatorial current sheet that originates at the tip of the closed line region, as first proposed in Paper II. The reason there exists an equatorial current sheet is that magnetic field lines leave one pole of the star, open up to infinity, and return from infinity to the other pole of the star. In doing so, they also carry the same amount of poloidal electric current in each hemisphere from each pole of the star to infinity. These two electric currents return to the star through the equatorial current sheet. In other words, the only reason an equatorial current sheet develops is simply to satisfy closure of the global poloidal electric current circuit (Contopoulos et al., 1999). It is thus obvious that, if we artificially (numerically) invert the direction of the field lines that leave the star from the southern pole, the electric current direction in the southern hemisphere will be inverted, and therefore there will be no need to close the global poloidal electric current circuit through an equatorial current sheet. This configuration is clearly artificial (it is equivalent to a magnetic monopole), but it is mathematically and dynamically equivalent to the configuration that we are investigating in the open line region, only without the mathematical discontinuity of the equatorial current sheet! We are able to implement this trick because we are treating the open line region independently from the closed line region. This is the same trick assumed by Bogovalov (1999) when he obtained the solution for the tilted split monopole. We here generalize his approach and show that it is also valid (and very helpful) in the numerical treatment of the open line region in the more general dipole magnetosphere. We have thus found a way to make the equatorial current sheet discontinuity ‘numerically disappear’. The ondulating surface separating open magnetic field lines that originate in the north stellar hemisphere from those that originate in the south is where the equatorial current sheet truly lies when the magnetic field direction in the southern stellar hemisphere is reversed back to its true value.
@@
Figure 1: Evolution of the total NN losses in the closed (blue) and open-line (orange) regions with respect to the number of training epochs (the difference in the number of epochs has to do with the different number of steps the training algorithm takes during its second-order optimization). We start the training with a first order optimizer and switch to a second order one after $10^{4}$ training epochs. For the first $10^{5}$ training epochs we kept the initial dipolar shape of the separatrix. Beyond that, we adjusted the shape of the separatrix every about $10^{4}$ epochs. Top: evolution during separatrix re-adjustment. Bottom: evolution during one final training for the fixed final shape of the separatrix. The achieved training loss between $10^{-4}$ and $10^{-5}$ is deemed satisfactory.
@@
Figure 2: Cross section of steady-state solution representing an inclined rotator with $\lambda=20^{\circ}$ and $\theta_{\rm pc}=36^{\circ}$. In this simulation $r_{*}=0.25R_{\rm lc}$. Rotation axis along $z$. The inclined magnetic axis lies along the corotating $xz$ plane shown. Thick black lines: separatrix between open and closed field lines. Red lines: the initial dipolar shape of the separatrix before readjustment. For this particular choice of the polar cap, the dipole is significantly stretched outwards closer to the light cylinder (represented by the two dashed lines at $x/R_{\rm lc}=\pm 1$). The equatorial current sheet lies where lines from the north and south polar cap rim meet. Color scale: ratio $B_{p}/B$. This represents the development of the azimuthal magnetic field $B_{\phi}$ accross the magnetosphere. Notice that at the magnetospheric Y-point where the equatorial current sheet connects to the separatrix current sheet, $B_{p}=0$ and $B_{\phi}\neq 0$ as expected (Uzdensky, 2003).
Figure 3: Same as Fig. 2 but along the corotating $yz$ plane.
@@
Figure 4: The steady-state solution of Fig. 2 in 3D as seen from above. Transparent yellow surface: light cylinder. We plot only open magnetic field lines from the rim of the northern polar cap. We see the clear azimuthal break of open field lines expected very close to the Y-point where $B_{p}\rightarrow 0$ and $B_{\phi}\neq 0$. This interesting feature was never observed so clearly before in previous numerical solutions.
@@
Figure 5: The steady-state solution of Fig. 4 as seen from the side. The ondulating shape of the equatorial current sheet is clearly seen. Notice that it does not develop the instabilities seen in all previous numerical solutions.
@@
Figure 7: Distribution of current parameter $\alpha$ along the stellar surface as seen from above the axis of rotation. $x/y$ axes along $\phi=0^{\circ}/90^{\circ}$ respectively. $\alpha=0$ in the green closed-line region outside the polar cap. Blue region: main magnetospheric current. Yellow-red region: part of the return current (the main part of the return current lies along the separatrix-black dotted line).
Figure 8: Evolution with distance of the total Poynting flux $L$ calculated over spheres of radius $r$ centered over the central star, normalized with respect to the aligned rotator’s canonical luminosity value $L_{0}\equiv B_{*}^{2}r_{*}^{6}c/(4R_{\rm lc}^{4})$. The value of $L$ agrees with previous estimates in the literature (dashed line according to eq. 7). Energy is (almost) conserved beyond the Y-point and the light cylinder due to the absence of dissipation in the magnetospheric current sheets.
@@ 3 Results
We experimented with various NN configurations, number layers and nodes per layer, various optimization procedures, and various activation functions. This tuning of hyperparameters is a tedious manual process that requires significant effort before the first meaningful results are obtained. In the solution presented below we implemented two Fully Connected NNs with 3 entires ($x,y,z$), 5 internal layers consisting of 128 nodes each, and 4 exits ($B_{x},B_{y},B_{z},\alpha$), first order Adam optimizers initially and second order Broyden-Fletcher-Goldfarb-Shanno (BFGS) optimizers later on, tanh activation functions, and roughly 200,000 training points distributed randomly over the various magnetospheric computational domains and their boundaries. Notice that, in order to avoid overfitting, the training points are periodically updated. The reason we implemented a cartesian system of coordinates was to avoid the singularities of the spherical system of coordinates around the axis $\theta=0$ and the requirement for periodicity between $\phi=0$ and $\phi=2\pi$. We validated our procedure by training our NNs with the analytic vacuum dipole solution of ×$\nabla\times{\bf B}=0$ (no rotation, no light cylinder, no electric currents). The comparison with the analytic dipolar solution was fair (see also Paper I). This gave us confidence to address the full 3D pulsar magnetosphere problem with PINNs. The separatrix is initially chosen to have a fixed dipolar shape for about 100,000 training epochs. After the initial training, the shape of the separatrix is re-adjusted every about 10,000 training epochs. In the top part of Fig. 1 we show the evolution of the losses for our main NNs during the re-adjustment of the shape of the separatrix. We performed 28 readjustments of the shape of the separatrix. We observed that after about 10 readjustments, the solution relaxes to a particular separatrix shape that satisfies pressure balance at all points accross it. In the bottom part of Fig. 1 we show the evolution of the losses during one final NN training for the fixed final shape of the separatrix (the observed spikes are due to changes in the training points). In total, the training of our two NNs required about $6\times 10^{5}$ training epochs and about 24 hours of computational time on a standard off-the-shelf computer with a GPU in order to reach a satisfactory solution with individual requirements satisfied to a precision of a few times $10^{-5}$ in both NNs. We thus obtained the first dissipationless ideal steady-state solution for a star with radius $r_{*}=0.25R_{\rm lc}$, a uniform polar cap with constant angular opening $\theta_{\rm pc}=36^{\circ}$ ($\pi/5$) around the magnetic axis, and an inclination angle of $\lambda=20^{\circ}$ ($\pi/9$). As we will see next, our first results are promising and emphasize the potential of our method.
@@
In Figs. 2 and 3 we show two cross-sections of the 3D solution along the corotating $xz$ plane that contains both the rotation and magnetic axes, and along the corotating $yz$ plane perpendicular to it. We observe that the initially dipolar separatrix is significantly stretched closer to the light cylinder. We also observe a clear nulling of the poloidal field $B_{p}$ at the Y-point where $B_{\phi}$ is non-zero due to the return current that flows back to the star along the equatorial current sheet. This result was predicted by Uzdensky (2003) but has never been seen before so clearly, mainly because no previous FFE solution obtained zero-thickness equatorial and separatrix current sheets. Uzdensky (2003) also predicted that, because $B_{p}=0$ and $B_{\phi}\neq 0$ right outside the Y-point, while $B_{\phi}\approx 0$ right inside, in order to achieve pressure balance between the inside and the outside at that point, $B_{p}$ must be non-zero just inside the Y-point. Therefore, technically, the shape of the Y-point must be a T as is clearly observed in Figs. 2 and 3. This result too has never been seen before in all previous numerical solutions, starting from Contopoulos et al. (1999), in which the separatrix is always shown with a pointed shape where the equatorial current sheet originates.
@@
In Figs. 4, 5 we show the 3D distribution of open magnetic field lines right above the magnetospheric equatorial and separatrix current sheet. Notice that the value of $\alpha$ is numerically adjusted along each open field line so that the latter smoothly crosses the light cylinder. It is interesting that while Contopoulos et al. (1999) developed a complex numerical method to perform this adjustment, NNs are able to accomplish this automatically. In Fig. 5 we clearly observe the ondulating shape of the equatorial current sheet beyond the closed line region.
@@
In Fig. 8 we calculate the integral of the Poynting vector ${\bf E}\times{\bf B}/(4\pi)$ over spheres of radius $r$ centered around the central star. We see that, because of the absence of numerical current sheets, the total Poynting flux is (almost) conserved in the open line region of our calculation. All previous numerical solutions show a clear drop ot the integrated Poynting flux beyond the light cylinder due to numerical dissipation at the equatorial current sheet. The obtained value of the total electromagnetic luminosity $L$ is comparable to the estimate of Spitkovsky (2006), namely
$L=1\pm 0.04\times L_{0}(1+\sin^{2}\lambda)\ ,$ (7)
from $r\approx 0.35R_{\rm lc}$ to about $2R_{\rm lc}$. Here $L_{0}=B_{*}^{2}r_{*}^{6}c/(4R_{\rm lc}^{4})$ is the canonical aligned rotator’s luminosity value, and $\lambda=20^{\circ}$ is the pulsar inclination angle. We will compute a complete sequence of pulsar inclination angles, polar cap openings, and polar cap shapes in Paper IV. We will then learn whether the result of Paper II is reproduced in 3D, namely whether magnetospheric solutions exist even for arbitrarily small polar caps with luminosities $L$ much smaller than the canonical value of eq. (7).

### v3

@@ Abstract
Results. In this paper we present our first results for an inclined dipole rotator using the new methodology. We are able to zoom-in around the Y-point and inside the closed-line region, and we observe new interesting features. This is the first time the steady-state 3D problem is addressed directly, and not through a time-dependent simulation that eventually relaxes to a steady-state.
Conclusions. We have trained a Neural Network that instantaneously yields the three components of the magnetic field and their spatial derivatives at any given point. Our results demonstrate the potential of the new method to generate new solutions of the ideal pulsar magnetosphere.
@@ 1 Gaps in our current understanding
Right after the discovery of pulsars (Hewish, 1968), Goldreich and Julian formulated the mathematical description of their plasma-filled magnetospheres in the ideal force-free limit in steady-state and axisymmetry (Goldreich & Julian, 1969). It took us three decades to produce the first solution of this idealized problem, mainly because of the mathematical singularity along the light cylinder (Contopoulos et al., 1999). Since then, this solution became a reference solution against which the results of numerical simulations are evaluated. As we will see below, the ideal force-free steady-state 3D problem is also well formulated mathematically, but still lacks a reference solution. This is mainly due to the presence of electric current sheets within the magnetosphere which strongly complicate all our standard numerical approaches.
@@
Recent advances in global particle-in-cell (PIC) simulations (e.g. Philippov et al., 2015; Hu & Beloborodov, 2022; Hakobyan et al., 2023; Bransgrove et al., 2023; Soudais et al., 2024) yield separatrix regions between open and closed field lines that exhibit a significant thickness beyond the simulation resolution (see Appendix), and a multi-layered internal structure. This is intriguing because, the same simulations are able to generate Harris-type equatorial current sheets with thickness at the limit of their numerical resolution (the skin-depth). Moreover, their closed-line regions terminate well inside the light cylinder, with regions of strong electromagnetic dissipation beyond their tips, something that was not expected by Goldreich and Julian. Interestingly, the shape of the tips resembles a pointed Y, thus deviating from the theoretically predicted T shape (Uzdensky, 2003). Furthermore, the extent of the closed-line region and the sharpness of the Y appears to be linked to the efficiency of pair production in the simulation (see comment in Kalapotharakos et al., 2018). Finally, our recent work in Contopoulos et al. (2024) indicates that as the closed-line region approaches the light cylinder, the electric current in the separatrix decreases, and in the limit where it touches the light cylinder, the equatorial current sheet vanishes entirely.
@@
The origin of these results remains unclear. Some researchers suspect that they arise either due to insufficient resolution in current PIC simulations, or to inadequate running times to reach a true steady-state. On the other hand, the resolution of current PIC simulations falls significantly short in modeling the microphysics of the equatorial current sheet. To generate pulsar light curves and spectra for comparison with observations, simulation results are extrapolated by several orders of magnitude (particle Lorentz factors, magnetic field values, electromagnetic spectra, etc.). Unfortunately, there is no consensus among research groups regarding the specific extrapolation method. Consequently, a universally accepted understanding of the physical origin of high-energy radiation from pulsars remains elusive.
@@
Our proposal is to revisit fundamental principles and independently derive the ideal force-free magnetosphere using a novel numerical approach that treats the separatrix and equatorial discontinuities as genuine ideal current sheets. In the present paper, we generalize in 3D the axisymmetric ideal solutions of Dimitropoulos et al. (2024, hereafter Paper I) and Contopoulos et al. (2024, hereafter Paper II). This is the third paper (Paper III) of an ongoing investigation that aims to produce the reference ideal solution of the 3D pulsar magnetosphere.
@@ 2 Domain decomposition
Once the size and shape of the polar cap is determined, the shape of the separatrix surface is unknown and must be self-consistently determined. In the present paper, we implement in 3D the procedure first proposed in Paper I. We begin by separating the regions of open and closed field lines with an ad hoc chosen surface, namely
@@
(see also Lyubarskii 1990). Treating such discontinuities inside any computational domain is very problematic and all computational methods generate spurious Gibbs oscillations around the separatrix (see discussion in Cao & Yang, 2022). We propose here a new way to treat such discontinuities as follows:
@@
We have implemented one more trick that greatly simplifies our problem, namely a novel numerical treatment of the equatorial current sheet that originates at the tip of the closed line region, as first proposed in Paper II. The reason there exists an equatorial current sheet is that magnetic field lines leave one pole of the star, open up to infinity, and return from infinity to the other pole of the star. In doing so, they also carry the same amount of poloidal electric current in each hemisphere from each pole of the star to infinity. These two electric currents return to the star through the equatorial current sheet. In other words, the only reason an equatorial current sheet develops is simply to satisfy closure of the global poloidal electric current circuit (Contopoulos et al., 1999). It is thus obvious that, if we artificially (numerically) invert the direction of the field lines that leave the star from the southern pole, the electric current direction in the southern hemisphere will be inverted, and therefore there will be no need to close the global poloidal electric current circuit through an equatorial current sheet. This configuration is clearly artificial (it is equivalent to a magnetic monopole), but it is mathematically and dynamically equivalent to the configuration that we are investigating in the open line region, only without the mathematical discontinuity of the equatorial current sheet! We are able to implement this trick because we are treating the open line region independently from the closed line region. This is the same trick assumed by Bogovalov (1999) when he obtained the solution for the tilted split monopole. We here generalize his approach and show that it is also valid (and very helpful) in the numerical treatment of the open line region in the more general dipole magnetosphere. We have thus found a way to make the equatorial current sheet discontinuity ‘numerically disappear’. The undulating surface separating open magnetic field lines that originate in the north stellar hemisphere from those that originate in the south is where the equatorial current sheet truly lies when the magnetic field direction in the southern stellar hemisphere is reversed back to its true value.
@@
Figure 1: Evolution of the total NN losses for the solution with $\lambda=20^{\circ}$ and $\theta_{\rm pc}=36^{\circ}$ in the closed-line (blue) and open-line (orange) regions with respect to the number of training epochs (the difference in the number of epochs has to do with the different number of steps the training algorithm takes during its second-order optimization). We started the training with a first order optimizer and switched to a second order one after $10^{4}$ training epochs. Beyond that, we adjusted the shape of the separatrix every $3\times 10^{4}$ epochs. The achieved total training loss between $10^{-5}$ and $10^{-6}$ is deemed satisfactory.
@@
Figure 2: Cross section of steady-state solution representing an inclined rotator with $\lambda=20^{\circ}$ and $\theta_{\rm pc}=36^{\circ}$. In this simulation $r_{*}=0.25R_{\rm lc}$. Rotation axis along $z$. The inclined magnetic axis lies along the corotating $xz$ plane shown. This particular cross section represents phase 00 (or $\pi$) of the corresponding rotating time-dependent solution. Closed thick black lines: separatrix between open and closed field lines. Open thick black lines: separatrix between open field lines that originate from the north and south magnetic poles. This is where the equatorial current sheet lies. Red lines: initial dipolar shapes of the separatrices before readjustment. For this particular choice of the polar cap, the dipole is significantly stretched outwards closer to the light cylinder (represented by the two dashed lines at $x/R_{\rm lc}=\pm 1$). The equatorial current sheet lies where lines from the north and south polar cap rim meet. Color scale: ratio $B_{p}/B$. This represents the development of the azimuthal magnetic field $B_{\phi}$ accross the magnetosphere. Notice that at the magnetospheric Y-point where the equatorial current sheet connects to the separatrix current sheet, $B_{p}=0$ and $B_{\phi}\neq 0$ as expected (Uzdensky, 2003).
Figure 3: Same as Fig. 2 but along the corotating $yz$ plane. This particular cross section represents phase $\pi/2$ (or $23\pi/$ 2) of the corresponding rotating time-dependent solution.
@@
Figure 4: The steady-state solution of Fig. 2 in 3D as seen from above. Yellow cylinder: light cylinder. Pale undulating surface: equatorial current sheet. We plot only open magnetic field lines just inside the rim of the northern polar cap. We see the clear azimuthal break of open field lines expected very close to the Y-point where $B_{p}\rightarrow 0$ and $B_{\phi}\neq 0$.
@@
Figure 5: The steady-state solution of Fig. 4 as seen from the side. We plot open magnetic field lines just inside the rims of both polar caps.The undulating shape of the equatorial current sheet between them is clearly seen. Notice that it does not develop the physical instabilities seen in all previous numerical solutions because this physics is missing in our work (we obtain the solution without a current sheet and then reverse the direction of the field in the southern magnetic hemisphere).
@@
Figure 7: Distribution of current parameter $\alpha$ along the stellar surface as seen from above the polar cap for the solution of Fig. 2. $\alpha=0$ in the green closed-line region outside the polar caps. $\phi=0$ along the corotating $xz$ plane of Fig. 2 in the direction of inclination of the magnetic axis. Blue region: main magnetospheric current. Yellow-red region: part of the return current (the main part of the return current lies along the separatrix-thick black dotted line).
Figure 8: Evolution with distance of the total Poynting flux $L$ calculated over spheres of radius $r$ centered over the central star, normalized with respect to the aligned rotator’s canonical luminosity value $L_{0}\equiv B_{*}^{2}r_{*}^{6}c/(4R_{\rm lc}^{4})$. The value of $L$ is close to previous estimates in the literature (dashed line according to Spitkovsky (2006) for the pulsar inclination angle considered). Energy is (almost) conserved beyond the Y-point and the light cylinder.
@@ 3 Results
We experimented with various NN configurations, number layers and nodes per layer, various optimization procedures, and various activation functions. This tuning of hyperparameters is a tedious manual process that requires significant effort before the first meaningful results are obtained. In the solution presented below we implemented two Fully Connected NNs with 3 entires ($x,y,z$), 5 internal layers consisting of 128 nodes each, and 4 exits ($B_{x},B_{y},B_{z},\alpha$), first order Adam optimizers initially and second order Broyden-Fletcher-Goldfarb-Shanno (BFGS) optimizers later on, tanh activation functions, and roughly 270,000 training points distributed randomly over the various magnetospheric computational domains and their boundaries. Notice that, in order to avoid overfitting, the training points are periodically updated. The reason we implemented a cartesian system of coordinates was to avoid the singularities of the spherical system of coordinates around the axis $\theta=0$ and the requirement for periodicity between $\phi=0$ and $\phi=2\pi$. We validated our procedure by training our NNs with the analytic vacuum dipole solution of ×$\nabla\times{\bf B}=0$ (no rotation, no light cylinder, no electric currents). The comparison with the analytic dipolar solution was fair (see also Paper I). This gave us confidence to address the full 3D pulsar magnetosphere problem with PINNs. In Fig. 1 we show the evolution of the total training losses for our main NNs. The shape of the separatrix was re-adjusted every 30,000 training epochs. We performed 25 readjustments (high spikes in Fig. 1) in order to reach a solution with individual requirements satisfied to a precision of a few times $10^{-6}$ in both NNs. The solution relaxes to a particular separatrix shape that satisfies pressure balance at all points accross it. The training of our two NNs required about 24 hours of computational time on a standard off-the-shelf computer with a GPU. We thus obtained a dissipationless ideal steady-state solution for a star with radius $r_{*}=0.25R_{\rm lc}$, a uniform polar cap with constant angular opening $\theta_{\rm pc}=36^{\circ}(\pi/5)$ around the magnetic axis, and an inclination angle of $\lambda=20^{\circ}(\pi/9)$. As we will see next, our first results are promising and emphasize the potential of our method.
@@
In Figs. 2 and 3 we show two cross-sections of the 3D solutions along the corotating $xz$ plane that contains both the rotation and magnetic axes, and along the corotating $yz$ plane perpendicular to it. We observe that the initially dipolar separatrices are significantly stretched closer to the light cylinder. We also observe a clear nulling of the poloidal field $B_{p}$ at the Y-point where $B_{\phi}$ is non-zero due to the return current that flows back to the star along the equatorial current sheet. This result was predicted by Uzdensky (2003) but has never been seen before so clearly, mainly because no previous ideal force-free solution obtained zero-thickness equatorial and separatrix current sheets. Uzdensky (2003) also predicted that, because $B_{p}=0$ and $B_{\phi}\neq 0$ right outside the Y-point, while $B_{\phi}\approx 0$ right inside, in order to achieve pressure balance between the inside and the outside at that point, $B_{p}$ must be non-zero just inside the Y-point. Therefore, technically, the shape of the Y-point must be a T as is clearly observed in Figs. 2 and 3. This result too has never been seen before in all previous numerical solutions, starting from Contopoulos et al. (1999), in which the separatrix is always shown with a pointed shape where the equatorial current sheet originates.
@@
In Figs. 4, 5 we show the 3D distribution of open magnetic field lines right above and below the magnetospheric equatorial and separatrix current sheet. Notice that the value of $\alpha$ is numerically adjusted along each open field line so that the latter smoothly crosses the light cylinder. It is interesting that while Contopoulos et al. (1999) developed a complex numerical method to perform this adjustment, NNs are able to accomplish this automatically. In Fig. 5 we clearly observe the undulating shape of the equatorial current sheet beyond the closed line region. This is determined as the boundary between the regions of open field lines that originate from the north and south magnetic poles. Notice that the undulating black lines in Figs. 2 and 3 were obtained from the intersection of the 3D undulating current sheet of Figs. 4 and 5 with the $xz$ and $yz$ planes respectively.
@@
In Fig. 8 we calculate the integral of the Poynting vector ${\bf E}\times{\bf B}/(4\pi)$ over spheres of radius $r$ centered around the central star. We see that, because of the absence of numerical current sheets, the total Poynting flux is (almost) conserved in the open line region of our calculation, and we observe a slight numerical dissipation of around $10\%$ up to about $3R_{\rm lc}$. All previous numerical solutions show a clear drop of the integrated Poynting flux beyond the light cylinder due to numerical dissipation at the equatorial current sheet. The obtained value of the total electromagnetic luminosity $L$ is comparable to the estimate of Spitkovsky (2006), namely
$L=1.2\pm 0.05\times L_{0}(1+\sin^{2}\lambda)\ ,$ (7)
from $r\approx 0.4R_{\rm lc}$ to about $3R_{\rm lc}$. Here $L_{0}=B_{*}^{2}r_{*}^{6}c/(4R_{\rm lc}^{4})$ is the canonical aligned rotator’s luminosity value, and $\lambda=20^{\circ}$ is the pulsar inclination angle. We will compute a complete sequence of pulsar inclination angles, polar cap openings, and polar cap shapes in Paper IV. We will then learn whether the result of Paper II is reproduced in 3D, namely whether magnetospheric solutions exist even for arbitrarily small polar caps with luminosities $L$ much smaller than the canonical value of eq. (7).